OpenStudy (anonymous):

a raindrop with a mass of 1 mg falls at 3 m/s when there is no wind. find its kinetic energy when a wind of 5 m/s is blowing.

OpenStudy (anonymous):

Find it's kinetic energy using the kinetic energy formula for 3 m/s and then for 2 m/ s since that's the difference between 3 and 5 and then add them up.

OpenStudy (rajat97):

sorry but @mihirb is wrong you may find the vector sum of the velocities which comes out to be (34)^1/2 and then substitute it in the formula k.e.=(mv^2)/2 so you may get the answer=10^-6 x 34 / 2 = 1.7 x 10^-5 J

OpenStudy (anonymous):

K.E. = (1/2)mv^2 Assuming the wind is horizonal to the ground, v = Sqrt( Vx^2 + Vy^2 ) = Sqrt(5^2 + 3^2) = Sqrt(34) = 5.83 m/s mass = 1 milligram = 1.0*10^(-3) gr = 1.0*10^(-6) kg K.E. = (1/2)*(1.0*10^(-6))*(5.83)^2 = 17 x10(-6) (kg m/s^2)